a particle moving with uniform retardation covers three successive equal distance. the average velocity during first and third parts of the journey being 20 meter/second and 12m/s respectively. determine the average velocity in the middle part of the journey.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
avg. vel. for first part : = x/t1 {total displacement/total time} = 20
avg. vel. for third part: = x/t3 = 12
avg. vel. for 2nd part = x/t2 = 12t3/t2 or = 20t1/t2